At what rate must the potential difference between the plates of a parallel-plate capacitor with a 2.0uF capacitance velocity changed to produce a displacement current of 1.5A
1
Expert's answer
2020-02-10T09:24:43-0500
"I=C\frac{dU}{dt}"

"\frac{dU}{dt}=\frac{I}{C}"

"\frac{dU}{dt}=\frac{1.5}{2\cdot 10^{-6}}=7.5\cdot 10^5\frac{V}{s}"